# Gradient Inversion Attack on Graph Neural Networks

### v1

@@ Appendix F Additional Graph Classification results on TUDatasets
For the graph classification task, we also conduct the attack on the following three datasets from the TUDataset repository Morris et al. (2020).
@@
  • MUTAG Debnath et al. (1991): It consists of $188$ chemical compounds from two classes. The vertices represent atoms and edges represent chemical bonds. Node features denote the atom type represented by one-hot encoding. Labels represent their mutagenic effect on a specific gram-negative bacterium.
@@
  • COIL-RAG Riesen & Bunke (2008): Dataset in computer vision where images of objects are represented as region adjacency graphs. Each node in the graph is a specific part in the original image. Graph labels denote the object type.
@@
  • FRANKENSTEIN Kazius et al. (2005): It contains graphs representing chemical molecules, where the vertices are chemical atoms and edges represent the bonding. Node features denote the atom type. Binary labels represent mutagenecity.
@@
The results are obtained by randomly selecting $20$ graphs from the datasets. The selected graphs are fixed for different parameter selections for fair comparison. In this case since the graphs are small we can directly use the Mean Absolute Error (MAE) of the recovered adjacency matrix to measure the performance of recovering the underlying graph structure. Specifically, the MAE is given as follows
@@
Also, instead of directly sampling edges from the probabilistic adjacency matrix as given in Section 3.2, the MAE is also evaluated with and without thresholding the recovered adjacency matrix. We apply min-max normalization to the resulting $\hat{\mathbf{A}}$ before thresholding. After applying the threshold, $\hat{\mathbf{A}}_{\tau}:=th(\hat{\mathbf{A}},\tau)$ is obtained, where $th(\mathbf{Z},\tau)$ is a entry-wise operator that returns 00 if $z_{ij}<\tau$, and returns $1$ otherwise.
@@
Table 12: Recovery of $\mathbf{A}$ (MAE) for Graph Attacker-a
$\tau$ MUTAG COIL-RAG FRANK.(×$\times 10^{-4}$)
@@
F.1 Graph Attacker-a
Table 5(d) shows the MAE of recovering $\mathbf{A}$ with thresholding (for different values of $\tau$) and without (N/A in the table). The attack achieves the best results on the FRANKENSTEIN dataset while using GraphSAGE. Each node in the FRANKENSTEIN dataset has $D=780$ features, whereas MUTAG and COIL-RAG have only $D=7$ and and $D=64$ features respectively. Moreover, the average number of nodes ($N$) per graph in FRANKENSTEIN, MUTAG and COIL-RAG are $16.9$, $17.9$ and $3.01$ respectively. This implies that $\mathbf{X}\in\mathbb{R}^{N\times D}$ is full-row rank with high probablity in COIL-RAG and FRANKENSTEIN. Meanwhile, the node features in the MUTAG dataset consists of only binary values. It can be observed from the experimental results that the best and worst results are obtained on FRANKENSTEIN and MUTAG respectively. This is also consistent with Proposition 6 which requires $\mathbf{X}$ to be full row-rank for successfully recovering $\tilde{\mathbf{A}}$. The recovery performance improves for COIL-RAG with GraphSAGE at $\tau=0.8$. The FRANKENSTEIN dataset is recovered accurately even with the GCN framework.
F.2 Graph Attacker-b
Table 13 lists the RNMSE of recovering $\mathbf{X}$. It can be observed that the attack is able to recover $\mathbf{X}$ with high accuracy for GraphSAGE across all three datasets. For the GCN framework, it can be observed that the algorithm fails to recover $\mathbf{X}$.
Table 13: Recovery of $\mathbf{X}$ (RNMSE) for Graph Attacker-b.
@@ F.2 Graph Attacker-b
Figure 6: Recovery of different graph structures (MAE) for Graph Attacker-c using GraphSAGE.
Table 14: Recovery of $\mathbf{A}$ (MAE) in for Graph Attacker-c with different thresholds.
@@
F.3 Graph Attacker-c
In this scenario, since the attacker has no knowledge of the graph data, we try different settings and observe their effects on the recovery.
The effect of initialization: As already been discovered by previous studies Gupta et al. (2022); Geiping et al. (2020), the quality of recovery to a large degree can depend on the initialization. The closer the initialization is to the original data, the better the recovery. An initialization value of $0.1$ implies that all the entries of $\hat{\mathbf{X}}$ and $\hat{\mathbf{A}}$ are initialized to $0.1$ and similarly $\mathcal{N}(0,1)$ implies that all entries are initialized from the standard normal distribution. Since the values of $\mathbf{X}$ and $\mathbf{A}$ lie in the range $[0,1]$ for all three datasets, we tested the effect of different initialization points in the same range. Specifically, we tested the initialization values as $\{0,0.1,0.5,1,\mathcal{N}(0,1)\}$. The results of the attack with different initialization strategies is shown in Table 16. The leftmost column gives the initialization values of the dummy adjacency and node feature matrices. For the adjacency matrix, the results are listed with and without thresholding. Similar to Scenario 2, we apply min-max normalization (before thresholding) and a threshold value of $0.5$ is chosen. A more compact version of the results showing the best performance over all initialization strategies is shown in Table 15. It can be observed that for GraphSAGE, the attack can accurately recover both $\mathbf{X}$ and $\mathbf{A}$ in the MUTAG and FRANKENSTEIN datasets. While using GCN, even though $\mathbf{X}$ cannot be recovered, $\mathbf{A}$ can be recovered accurately.
For GraphSAGE in Table LABEL:result:table_init_all_sage, it can be observed that even without any knowledge of the graph data, the attack can accurately recover both $\mathbf{X}$ and $\mathbf{A}$ just by utilising a good initialization strategy. For example, in Table LABEL:result:table_init_all_sage for the FRANKENSTEIN dataset, the attack gives a high error for both $\mathbf{A}$ and $\mathbf{X}$ while using $\mathcal{N}(0,1)$ as the initialization. However, the MAE drops by orders of magnitude when we initialize $\mathbf{X}$ and $\mathbf{A}$ to a value of $0.2$. Even for the GCN framework in Table LABEL:result:table_init_all_gcn, while the data cannot be recovered in most of the cases, it can be observed that a good initialization point can leak $\mathbf{A}$ in COIL-RAG and FRANKENSTEIN datasets.
Table 15: Recovery of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (MAE) for Graph Attacker-c.
@@ F.3 Graph Attacker-c
The effect of threshold: To explore the effect of thresholding on the adjacency matrix, we examined the MAE across all three datasets using different threshold values. Table 6(e) lists the MAE of recovering $\mathbf{A}$ at different thresholds. It can be observed that unlike Scenario 2, the performance of the algorithm stays roughly the same for both GraphSAGE and GCN. Also, the FRANKENSTEIN dataset is no longer vulnerable to the attack unlike Scenario 2. Few of the threshold values reduce the MAE by a considerable amount. For example, $\tau=0.8$ works well for MUTAG and FRANKENSTEIN in the case of GraphSAGE, and for MUTAG in the case of GCN.
Vulnerable graph structures: In order to identify vulnerable graph structures, we tested the performance of the attack in attacking graphs with varying number of nodes and edges for the GraphSAGE framework. Specifically, the performance was tested on Erdos-Renyi graphs (with varying edge probabilities). Figure 6 shows the MAE of recovering $\mathbf{A}$ for different graphs. The number of features for all graphs is set to $D=10$, and their values are sampled from the standard normal distribution. Each subfigure represents different initialization strategies for $\hat{\mathbf{X}}$ and $\hat{\mathbf{A}}$. For each graph with a specific number of nodes, the MAE is averaged over $20$ runs of the attack. As the number of nodes in a graph increases, the quality of recovery decreases. More importantly, it is evident that the vulnerability to the attack increases with decreasing sparsity. Fully connected graphs (ER graphs with edge probability $1.0$) are the least vulnerable to the attack .
Table 16: Recovery of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (MAE) for Graph Attacker-c .

### v2

@@ Appendix G Additional Graph Classification results on TUDatasets
For the graph classification task, we conduct the attack on the following three datasets from the TUDataset repository (Morris et al., 2020).
@@
  • MUTAG (Debnath et al., 1991): It consists of 188188188 chemical compounds from two classes. The vertices represent atoms and edges represent chemical bonds. Node features denote the atom type represented by one-hot encoding. Labels represent their mutagenic effect on a specific gram-negative bacterium.
@@
  • COIL-RAG (Riesen & Bunke, 2008): Dataset in computer vision where images of objects are represented as region adjacency graphs. Each node in the graph is a specific part in the original image. Graph labels denote the object type.
@@
  • FRANKENSTEIN (Kazius et al., 2005): It contains graphs representing chemical molecules, where the vertices are chemical atoms and edges represent the bonding. Node features denote the atom type. Binary labels represent mutagenecity.
The average number of nodes in the graphs of these datasets is significantly smaller than social networks (see Table 13). However, unlike social networks, these graphs might not necessarily exhibit properties like feature smoothness and sparsity. Hence, for all the attacks we use the objective function as defined in equation 12. The results in all attack settings are again consistent with the theoretical analysis.
@@
The results are obtained by randomly selecting 202020 graphs from the datasets. The selected graphs are fixed for different parameter selections for fair comparison. In this case since the graphs are small we can directly use the Mean Absolute Error (MAE) of the reconstructed adjacency matrix to measure the performance of reconstructing the underlying graph structure. Specifically, the MAE is given as follows
@@
Also, instead of directly sampling edges from the probabilistic adjacency matrix as given in Section 4.2, the MAE is also evaluated with and without thresholding the reconstructed adjacency matrix. We apply min-max normalization to the resulting $\hat{\mathbf{A}}$ before thresholding. After applying the threshold, $\hat{\mathbf{A}}_{\tau}:=th(\hat{\mathbf{A}},\tau)$ is obtained, where $th(\mathbf{Z},\tau)$ is a entry-wise operator that returns 0 if $z_{ij}<\tau$, and returns 111 otherwise.
@@
Table 14: Reconstruction of $\mathbf{A}$ (MAE) for Graph Attacker-G
[…]
@@
Graph Attacker-G.
Table 5(d) shows the MAE of reconstructing $\mathbf{A}$ with thresholding (for different values of $\tau$) and without (N/A in the table). The attack achieves the best results on the FRANKENSTEIN dataset while using GraphSAGE. Each node in the FRANKENSTEIN dataset has $D=780$ features, whereas MUTAG and COIL-RAG have only $D=7$ and $D=64$ features respectively. Moreover, the average number of nodes ($N$) per graph in FRANKENSTEIN, MUTAG and COIL-RAG are 16.916.916.9, 17.917.917.9 and 3.013.013.01 respectively. This implies that $\mathbf{X}\in\mathbb{R}^{N\times D}$ is full-row rank with high probability in COIL-RAG and FRANKENSTEIN. Meanwhile, the node features in the MUTAG dataset consist of only binary values. It can be observed from the experimental results that the best and worst results are obtained on FRANKENSTEIN and MUTAG respectively. This is also consistent with Proposition 6 which requires $\mathbf{X}$ to be full row-rank for successfully reconstructing $\tilde{\mathbf{A}}$. The reconstruction performance improves for COIL-RAG with GraphSAGE at $\tau=0.8$. The FRANKENSTEIN dataset is reconstructed accurately even with the GCN framework.
Graph Attacker-N.
Table 15 lists the RNMSE of reconstructing $\mathbf{X}$. It can be observed that the attack is able to reconstruct $\mathbf{X}$ with high accuracy for GraphSAGE across all three datasets. For the GCN framework, it can be observed that the algorithm fails to reconstruct $\mathbf{X}$.
Table 15: Reconstruction of $\mathbf{X}$ (RNMSE) for Graph Attacker-N.
@@
Figure 6: Reconstruction of different graph structures (MAE) for Graph Attacker-GN using GraphSAGE.
Table 16: Reconstruction of $\mathbf{A}$ (MAE) in for Graph Attacker-GN with different thresholds.
@@
Graph Attacker-GN.
In this scenario, since the attacker has no knowledge of the graph data, we try different settings and observe their effects on the reconstruction. To improve the attack for Graph Attacker-GN where we have no prior knowledge of the graph, we utilize various initialization and thresholding strategies. Specifically, since the optimization process given in the algorithms can be greatly improved by utilizing different starting points, we initialize the dummy node feature and adjacency matrices with different values. In the case of thresholding, instead of sampling the edges given the probabilistic adjacency matrix $\hat{\mathbf{A}}^{P+1}$, we use various threshold values to sample edges.
The effect of initialization: Previous studies have established that the quality of reconstruction heavily depends on the initialization (Gupta et al., 2022; Geiping et al., 2020). The closer the initialization is to the original data, the better the reconstruction. An initialization value of 0.10.10.1 implies that all the entries of $\hat{\mathbf{X}}$ and $\hat{\mathbf{A}}$ are initialized to 0.10.10.1 and similarly $\mathcal{N}(0,1)$ implies that all entries are initialized from the standard normal distribution. Since the values of $\mathbf{X}$ and $\mathbf{A}$ lie in the range [0,1][0,1][ 0 , 1 ] for all three datasets, we tested the effect of different initialization points in the same range. Specifically, we tested the initialization values as $\{0,0.1,0.5,1,\mathcal{N}(0,1)\}$. The results of the attack with different initialization strategies is shown in Table 18. The leftmost column gives the initialization values of the dummy adjacency and node feature matrices. For the adjacency matrix, the results are listed with and without thresholding. Similar to Scenario 2, we apply min-max normalization (before thresholding) and a threshold value of 0.50.50.5 is chosen. A more compact version of the results showing the best performance over all initialization strategies is shown in Table 17. It can be observed that for GraphSAGE, the attack can accurately reconstruct both $\mathbf{X}$ and $\mathbf{A}$ in the MUTAG and FRANKENSTEIN datasets. While using GCN, even though $\mathbf{X}$ cannot be recovered, $\mathbf{A}$ can be reconstructed accurately.
For GraphSAGE in Table LABEL:result:table_init_all_sage, it can be observed that even without any knowledge of the graph data, the attack can accurately reconstruct both $\mathbf{X}$ and $\mathbf{A}$ just by utilising a good initialization strategy. For example, in Table LABEL:result:table_init_all_sage for the FRANKENSTEIN dataset, the attack gives a high error for both $\mathbf{A}$ and $\mathbf{X}$ while using $\mathcal{N}(0,1)$ as the initialization. However, the MAE drops by orders of magnitude when we initialize $\mathbf{X}$ and $\mathbf{A}$ to a value of 0.20.20.2. Even for the GCN framework in Table LABEL:result:table_init_all_gcn, while the data cannot be reconstructed in most of the cases, it can be observed that a good initialization point can leak $\mathbf{A}$ in COIL-RAG and FRANKENSTEIN datasets.
Table 17: Reconstruction of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (MAE) for Graph Attacker-GN.
@@
The effect of threshold: To explore the effect of thresholding on the adjacency matrix, we examined the MAE across all three datasets using different threshold values. Table 6(e) lists the MAE of reconstructing $\mathbf{A}$ at different thresholds. It can be observed that unlike Scenario 2, the performance of the algorithm stays roughly the same for both GraphSAGE and GCN. Also, the FRANKENSTEIN dataset is no longer vulnerable to the attack unlike Scenario 2. Few of the threshold values reduce the MAE by a considerable amount. For example, $\tau=0.8$ works well for MUTAG and FRANKENSTEIN in the case of GraphSAGE, and for MUTAG in the case of GCN.
Table 18: Reconstruction of $\mathbf{X}$ (RNMSE) and $\mathbf{A}$ (MAE) for Graph Attacker-GN.
@@
Vulnerable graph structures: In order to identify vulnerable graph structures, we tested the performance of the attack in attacking graphs with varying numbers of nodes and edges for the GraphSAGE framework. Specifically, the performance was tested on Erdos-Renyi graphs (with varying edge probabilities). Figure 6 shows the MAE of reconstructing $\mathbf{A}$ for different graphs. The number of features for all graphs is set to $D=10$, and their values are sampled from the standard normal distribution. Each subfigure represents different initialization strategies for $\hat{\mathbf{X}}$ and $\hat{\mathbf{A}}$. For each graph with a specific number of nodes, the MAE is averaged over 202020 runs of the attack. As the number of nodes in a graph increases, the quality of reconstruction decreases. More importantly, it is evident that the vulnerability to the attack increases with decreasing sparsity. Fully connected graphs (ER graphs with edge probability 1.01.01.0) are the least vulnerable to the attack.
Table 19: Reconstruction performance (GSM-0, GSM-1, GSM-2, and Full) on Tox21 results using GCN.
GSM-0 GSM-1 GSM-2 FULL
GLG (Ours) $80.86^{+1.5}_{-1.2}$ $96.51^{+1.1}_{-3.3}$ $97.02^{+0.8}_{-1.1}$ $98.71^{+1.0}_{-2.8}$
GRAIN $86.9^{+4.2}_{-5.7}$ $83.9^{+5.2}_{-6.9}$ $82.6^{+5.7}_{-7.4}$ 68.0±1.768.0$\pm$ 1.768.0 ± 1.7
DLG $31.8^{+4.5}_{-4.3}$ $20.3^{+5.5}_{-4.8}$ $22.8^{+6.6}_{-5.6}$ 1.0±0.21.0$\pm$ 0.21.0 ± 0.2
DLG + A $54.7^{+3.9}_{-4.2}$ $60.1^{+4.6}_{-5.2}$ $76.7^{+3.6}_{-4.8}$ 1.0±0.21.0$\pm$ 0.21.0 ± 0.2
TabLeak $25.1^{+5.1}_{-4.3}$ $12.4^{+5.5}_{-4.3}$ $10.8^{+5.6}_{-3.9}$ 1.0±0.21.0$\pm$ 0.21.0 ± 0.2
TabLeak + A $55.6^{+3.9}_{-3.9}$ $57.7^{+4.1}_{-4.6}$ $73.8^{+2.8}_{-3.5}$ 1.0±0.21.0$\pm$ 0.21.0 ± 0.2
Table 20: Reconstruction of $\mathbf{X}$ (RNMSE $\downarrow$) and $\mathbf{A}$ (AUC $\uparrow$ and AP $\uparrow$) with Graph Attacker-GN on Tox21 dataset using GCN.
Framework $\mathbf{X}\downarrow$ AUC $\uparrow$ AP $\uparrow$
GLG (Ours) 0.76$\pm$0.13 0.93$\pm$0.10 0.94$\pm$0.04
GRAIN 1.35$\pm$0.03 0.91$\pm$0.10 0.65$\pm$0.11
Appendix H Comparison with Other Methods
In this section, we compare our proposed gradient inversion attack, GLG, against several existing methods, including the only other graph-specific gradient inversion approach, GRAIN (Drencheva et al., 2025), as well as two widely used baselines originally designed for image or tabular data, DLG and TabLeak. We evaluate our approach on the Tox21 dataset, a well-known chemical dataset from the MoleculeNet benchmark (Wu et al., 2018). The task is to reconstruct both the node features $\mathbf{X}$ and the adjacency matrix $\mathbf{A}$ from gradients alone.
[…]
H.1 Detailed Analysis
We evaluate our method under the Graph attacker-GN threat model, which attempts to reconstruct both the node feature matrix $\mathbf{X}$ and the adjacency matrix $\mathbf{A}$ given the gradients. Table 19 shows that while reconstruction of $\mathbf{X}$ remains challenging, as indicated by a relatively lower GSM-0 score, reconstruction of the graph structure (reflected in the GSM-1, GSM-2 and FULL scores) is significantly improved compared to other methods. Table 20 presents the RNMSE, AUC, and AP results for a more detailed comparison between GLG and GRAIN.
Observations.
  • Node Feature Reconstruction (GSM-0): The GSM-0 score measures the fidelity of the node feature reconstruction. Our experiments indicate that reconstructing $\mathbf{X}$ from gradients is challenging due to the lack of direct node-level supervision during backpropagation through graph convolutional layers.
[…]
  • Comparison with GRAIN: Although GRAIN is tailored for graph gradient inversion attacks, our method achieves higher scores in capturing global structure as indicated by improved GSM-1, GSM-2, and FULL metrics. Table 20 shows a similar trend: GLG more effectively reconstructs the graph structure, as indicated by higher AUC and AP, while also slightly improving nodal feature reconstruction, as reflected by the lower RNMSE.
  • Baseline Methods (DLG and TabLeak): These methods, originally developed for images and tabular data, underperform in the graph domain. Even when augmented with the adjacency matrix information (“+A”), their performance across all GSM metrics remains inferior, highlighting the unique challenges of gradient inversion attacks on graph data.
Interpretation.
The lower GSM-0 score suggests that reconstructing node features $\mathbf{X}$ purely from gradient information is non-trivial. However, the significant improvements observed in GSM-1, GSM-2, and FULL indicate that the model is highly effective in capturing the graph structure. This is critical, as many practical applications depend on the accurate reconstruction of the graph topology rather than on precise node-level details.
Conclusion.
[…]